can you find the following slopes with these numbers???
(3,4) and (7,5)

Answer by blwinbbbles(106)

The slope (m) of a line is defined by:
m = (Y2 - Y1)/(X2 - X1) so: with the following points (3,4) and (7,5)
m = (5-4)/(7-3) = (1)/(4) so m = 1/4
so now you know the slope is 1/4
Hope this helps